# Reconstructing the unitary part of a noisy quantum channel

### v2

@@ V Resource analysis
The so-called sampling complexity $t$ takes into account the total number of measurements (settings and repetitions) to obtain an estimate for a certain state, respectively channel. For example, the sampling complexity for state tomography of a full-rank density matrix is $t=\mathcal{O}(d^{4})=\mathcal{O}(16^{N})$ [18]. Compressed sensing allows one to reduce the sampling complexity $t$ for a state with rank $r$ to [19]
@@
In state-based reconstruction, it is the rank of the output states that matters, and we set $r\equiv r_{\text{out}}$ in the following with
@@
In general, the different input states of the mixed-state and pure-state reconstruction will all have a different rank on output under a non-unitary dynamical map. We cover all such circumstances by taking a worst-case estimate, i.e., $r_{\text{out}}$ is the maximum rank among all output states. Moreover, we need to account for the sampling complexity of all $n_{\text{out}}$ output states in Eq. (20). In this account, we deliberately ignore the contribution of the state $\rho_{P}$ to the resource scaling for two reasons: (i) It is needed for both the pure-state and mixed-state reconstruction, so when comparing the resource scaling of both, its contribution can be discarded. (ii) Since $\rho_{P}$ is a single pure state, independent of the system size $N$, its contribution to the overall resource scaling is vanishingly small. Under this assumption, $n^{\text{Mix}}_{\text{out}}=1$ for the mixed-state reconstruction, corresponding to the output state $\mathcal{D}(\rho_{B})$. In contrast, the pure-state reconstruction has $n^{\text{Pure}}_{\text{out}}=d$ output states, namely $\mathcal{D}(\rho_{B}^{(i)})$ with $i=1,\ldots,d$. Given $n_{\text{out}}$ and an estimate of the output states’ rank, the minimum number of channel uses the pure and mixed-state reconstruction $M$ is
@@
The rank of the mixed input state is $r_{\text{in}}=\text{rank}(\rho_{B})=d$. If the dynamical map is unitary, $\mathcal{D}(\cdot)\equiv\mathcal{D}_{U}(\cdot)$, then $r_{\text{out}}=\text{rank}(\mathcal{D}_{U}(\rho_{B}))=d$. If the dynamical map is non-unitary, $r_{\text{out}}=\text{rank}(\mathcal{D}(\rho_{B}))\leq d$. However, the mixed-state reconstruction breaks down if the output states are not full rank anymore, as shown in Fig. 1, since one can then not identify any longer the basis states by ordering the eigenvalues. Therefore we have to limit ourselves to $r_{\text{out}}=d$, which is as an additional weakness of the mixed-state reconstruction, compared to the pure-state reconstruction. With Eq. (20) and (22), the minimal number of channel uses for the mixed-state reconstruction becomes
@@
Here the superscript $(U)$ in $M_{\text{Mix}}^{(U)}$ symbolizes that the dynamics are unitary or sufficiently close to unitary such that the rank of the output states is not appreciably affected. The absence of this superscript indicates the general case of (potentially strong) non-unitary dynamics. Equation (23) demonstrates that the number of channel uses of the mixed-state reconstruction is not influenced by the unitarity of the underlying dynamical map, i.e, $M_{\text{Mix}}^{(U)}=M_{\text{Mix}}$. This is because the mixed initial state $\rho_{B}$ has full rank already and therefore, the output states have also full rank. Here we assume that the dynamical map is faithful in the sense that the full rank of the mixed input density matrix is preserved and no degeneracies occur. Otherwise, the use of the mixed-state reconstruction is prohibited anyway.
We now turn to the resource scaling of the pure-state reconstruction. The pure input states have rank $r^{(i)}_{\text{in}}=\text{rank}(\rho_{B}^{(i)})=1$. If the dynamical map is perfectly unitary $\mathcal{D}(\cdot)\equiv\mathcal{D}_{U}(\cdot)$ or sufficiently close to unitary $\mathcal{D}(\cdot)\approx\mathcal{D}_{U}(\cdot)$, then $r^{(i)}_{\text{out}}=\text{rank}(\mathcal{D}_{U}(\rho_{B}^{(i)}))=1$. We therefore get
@@
For the Choi reconstruction in Eq. (9), the minimum number of channel uses is $M=\mathcal{O}(N\hskip 2.84544pt16^{N})$ [17]. Under the assumption of a rank-1 Choi matrix using compressed sensing [18] and direct fidelity estimation [20], this reduces to [4]
@@
We conclude that the pure-state reconstruction requires the least amount of channel uses when the dynamics is unitary or close to unitary and is therefore the best choice in terms of required resources. (ii) The dissipative part strongly affects the rank of the states such that every output state has approximately full rank, $r_{\text{out}}=d$. In this case, Eq. (26) cannot be used since the rank-1 assumption is violated, and an estimate of the scaling for the Choi reconstruction is difficult [4, 19]. We will thus not compare with compressed sensing estimates of $M_{\text{Choi}}$ for this scenario in the following. A comparison of Eq. (23) with Eq. (25) shows
@@
For strong dissipation, the mixed-state reconstruction outperforms the pure-state reconstruction in terms of channel uses. Intuitively, this follows from the the fact that with stronger dissipation, the number of channel uses for the pure-state construction increases whereas it remains constant for the mixed-state reconstruction. Despite its worse scaling, there are three reasons why the pure-state reconstruction may still be the better choice also for strong dissipation: First, the gate error of the mixed-state reconstruction is higher than the pure-state one. Second, mixed states are experimentally much harder to initialize than pure states. Finally, the mixed-state reconstruction is more susceptible to degeneracies in the image, limiting the regime for which it can be used in practice.
@@
Finally, comparing Eq. (26) with the scaling of the pure-state reconstruction in Eq. (25), we assert that our method performs significantly better if $r_{\text{out}}\approx 1$ since the minimal number of channel uses becomes $M_{\text{Pure}}=\mathcal{O}(8^{N})\ll\mathcal{O}(16^{N})$. Even for more substantial dissipation, when $r_{\text{out}}\approx\sqrt{d}$, both methods perform equally well since $M_{\text{Pure}}=\mathcal{O}(2^{N}8^{N})=\mathcal{O}(16^{N})=M^{(U)}_{\text{Choi}}$. Hence, our method decreases the number of channel uses necessary to reconstruct the unitary part of a general dynamical map compared to previous work using compressed sensing, as long as $r_{\text{out}}\leq\sqrt{d}$, at the cost of larger reconstruction errors compared to Choi reconstruction. When comparing with Standard Quantum Process Tomography (SQPT), which requires $M_{\text{SQPT}}=\mathcal{O}(N16^{N})$ channel uses [17], the reduction in the required resources is even more substantial since $M^{(U)}_{\text{Choi}}<M_{\text{SQPT}}$.
@@ VI Conclusions
We have presented two methods to approximate the unitary part of a non-unitary evolution using a set of $d+1$ pure, resp. two mixed, input states, where $d$ is the dimension of Hilbert space. We have shown that reconstruction works well for no or weak decoherence. As expected, it struggles for strong decoherence where degeneracies in the image of the dynamical map hamper the meaningful identification of a specific unitary. We find that reconstruction using pure input states offers a good compromise between accuracy and resource cost. If accuracy is paramount, the Choi reconstruction provides the best results but requires a number of channel uses that quickly becomes unfeasible with increasing system size. When comparing to compressed sensing methods [18, 19, 4], in the regime of weak decoherence, the minimum number of channel uses is decreased from $M=\mathcal{O}(16^{N})$ to $M=\mathcal{O}(8^{N})$, where $N$ is the number of qubits, for the pure-state reconstruction. Our approach differs from tomographic approaches which rely on a least-squares fit over sets of quantum channel [4]. Although this precludes to obtain analytical bounds using techniques from, e.g., compressed sensing, the lack of a numerical optimization removes a point of failure and reduces computational effort.
@@ Appendix B CNOT-like gates for multiple qubits
For $N=2$, we consider the CNOT gate where the first (second) qubit is the control (target) qubit. For multiple qubits, i.e. $N\in\{3,4,5\}$, we assume that the extra qubits (as compared to the $N=2$ case) are additional control qubits while the target qubit remains the second one. We also assume that only a single qubit, namely the target qubit dissipates, so that the absolute dissipation is the same for every system size. Figures 6 and 7 show the gate error of the pure-state and mixed-state reconstruction methods respectively. They demonstrate that both methods can be used to reconstruct CNOT-like gates for multiple qubits, as indicated by the continuous gate error graphs. However, in the left panel of Fig. 7 we notice that for the mixed-state reconstruction there are degeneracies in the image (see red graph), that do not appear in the pure-state reconstruction, cf. left panel of Fig. 6. When comparing the gate errors of both methods, we again notice that the pure-state reconstruction performs better (compare position of the horizontal dotted blue line on the y-axis for the pure-state and mixed-state reconstruction).
Figure 6: Behavior of gate error $1-F_{G}(U_{0},U_{RC})$ of the pure-state reconstruction method of CNOT-like gates for four designated $T_{1}$ times as a function of the number of qubits NN.
Figure 7: Behavior of gate error $1-F_{G}(U_{0},U_{RC})$ of the mixed-state reconstruction method of CNOT-like gates for four designated $T_{1}$ times as a function of the number of qubits NN.

### v3

@@ V Resource analysis
The so-called sampling complexity $t$ accounts for the total number of measurements to obtain an estimate for a certain state, respectively channel. For example, the sampling complexity for state tomography of a full-rank density matrix is $t=\mathcal{O}(d^{4})=\mathcal{O}(16^{N})$ [21]. Compressed sensing allows one to reduce the sampling complexity $t$ of a density matrix with rank $r$ to [22]
@@
Since it is the rank of the output density matrix that matters for the reconstruction, we set $r\equiv r_{\text{out}}$. In general, the rank of each output state obtained from the non-unitary time evolution of the mixed, respectively pure initial states may be different. We cover all such circumstances by taking a worst-case estimate, i.e., $r_{\text{out}}$ is the maximum rank among all output states,
@@
Equation (21) captures the sampling complexity of a single output state which, for the overall resource estimate of the reconstruction, needs to be multiplied by the number of output states, $n_{\text{out}}$, required by the respective reconstruction method. In this respect, it is possible to ignore the contribution of the state $\rho_{P}$ to the resource scaling for several reasons: (i) $\rho_{P}$ is needed for both pure-state and mixed-state reconstruction, so when comparing the resource scaling of both, its contribution can be discarded. (ii) Since $\rho_{P}$ is a single pure state, independent of the system size $N$, its contribution to the overall resource scaling is comparatively small. (iii) There are no significant state preparation costs for $\rho_{P}$ as it involves only single qubit Hadamard gates on each qubit which in most modern architectures. Neglecting the contribution of $\rho_{P}$, $n^{\text{Mix}}_{\text{out}}=1$ for the mixed-state reconstruction, corresponding to the output state $\mathcal{D}(\rho_{B})$. In contrast, the pure-state reconstruction has $n^{\text{Pure}}_{\text{out}}=d$ output states, namely $\mathcal{D}(\rho_{B}^{(i)})$ with $i=1,\ldots,d$.
Given $n_{\text{out}}$ and an estimate of the output states’ rank, the minimum number $M$ of channel uses is
@@
We now evaluate this more concretely for mixed-state and pure-state reconstruction. While the rank of the mixed input state is $r_{\text{in}}=\text{rank}(\rho_{B})=d$, $r_{\text{out}}=\text{rank}(\mathcal{D}(\rho_{B}))\leq d$ where the inequality occurs for strongly non-unitary dynamical maps. Since, however, the mixed-state reconstruction breaks down if the output states are not full rank, as shown in Fig. 1, we only consider $r_{\text{out}}=d$ in the following. With Eqs. (21) and (23), the minimal number of channel uses for the mixed-state reconstruction then becomes
@@
Here, the superscript $(U)$ in $M_{\text{Mix}}^{(U)}$ symbolizes that the dynamics are unitary or sufficiently close to unitary such that the rank of the output states is not appreciably affected. The absence of this superscript indicates the general case of (potentially strong) non-unitary dynamics. We assume that the non-unitary dynamical map is faithful in the sense that the full rank of the mixed input density matrix is preserved and no degeneracies occur. In this case, the number of channel uses for the mixed-state reconstruction does not change for unitary or non-unitary dynamics, i.e. $M_{\text{Mix}}^{(U)}=M_{\text{Mix}}$.
We now turn to the resource scaling of the pure-state reconstruction. The pure input states have rank $r^{(i)}_{\text{in}}=\text{rank}(\rho_{B}^{(i)})=1$. If the dynamical map is perfectly unitary $\mathcal{D}(\cdot)\equiv\mathcal{D}_{U}(\cdot)$ or sufficiently close to unitary $\mathcal{D}(\cdot)\approx\mathcal{D}_{U}(\cdot)$, then $r^{(i)}_{\text{out}}=\text{rank}(\mathcal{D}_{U}(\rho_{B}^{(i)}))=1$. We therefore obtain
@@
For the Choi reconstruction in Eq. (9), the minimum number of channel uses is $M_{\text{Choi}}=\mathcal{O}(N\hskip 2.84544pt16^{N})$ [20]. The most efficient way to implement Choi reconstruction uses compressed sensing [21] and direct fidelity estimation [23]. Under the assumption of a rank-1 Choi matrix, the number of channel uses then reduces to [4]
@@
We conclude that the pure-state reconstruction requires the least amount of channel uses when the dynamics is unitary or close to unitary and is therefore the best choice in terms of required resources. (ii) The dissipative part strongly affects the rank of the output states such that every output state has approximately full rank, $r_{\text{out}}=d$. In this case, Eq. (27) cannot be used since the rank-1 assumption is violated, and an estimate of the scaling for the Choi reconstruction is difficult [4, 22]. We will thus not compare with compressed sensing estimates of $M_{\text{Choi}}$ for this scenario, and a comparison of Eq. (24) with Eq. (26) shows
@@
For strong dissipation, the mixed-state reconstruction outperforms the pure-state reconstruction in terms of channel uses. Intuitively, this follows from the fact that with stronger dissipation, the number of channel uses for the pure-state construction increases whereas it remains constant for the mixed-state reconstruction. Despite its worse scaling, there are three reasons why the pure-state reconstruction may still be the better choice also for strong dissipation: First, the reconstruction error of the mixed-state reconstruction is higher than the pure-state one. Second, mixed states are experimentally much harder to initialize than pure states. Finally, the mixed-state reconstruction is more susceptible to degeneracies in the image, limiting the regime for which it can be used in practice.
@@
Finally, comparing Eq. (27), representing the state of the art in channel reconstruction, with the scaling of the pure-state reconstruction in Eq. (26), we assert that our method performs significantly better if $r_{\text{out}}\approx 1$ since the minimal number of channel uses becomes $M_{\text{Pure}}=\mathcal{O}(8^{N})\ll\mathcal{O}(16^{N})$. Even for more substantial dissipation, when $r_{\text{out}}\approx\sqrt{d}$, both methods perform equally well since $M_{\text{Pure}}=\mathcal{O}(2^{N}8^{N})=\mathcal{O}(16^{N})=M^{(U)}_{\text{Choi}}$. Hence, our method decreases the number of channel uses necessary to reconstruct the unitary part of a general dynamical map compared to previous work using compressed sensing, as long as $r_{\text{out}}\leq\sqrt{d}$, at the cost of larger reconstruction errors compared to Choi reconstruction. When comparing with Standard Quantum Process Tomography (SQPT), which requires $M_{\text{SQPT}}=\mathcal{O}(N16^{N})$ channel uses [20], the reduction in the required resources is even more substantial since $M^{(U)}_{\text{Choi}}<M_{\text{SQPT}}$.
Figure 7: Minimum number of channel uses $M$ required for the pure-state and mixed-state reconstruction versus the number of qubits for different dissipation strengths. To avoid any bias due to the rank of the output states, $M$ is averaged over $n=\{100,75,50,10,10\}$ random unitaries for $N=\{2,3,4,5,6\}$ qubit assuming identical dissipation for each gate. Weak dissipation corresponds to $T_{2}=25T_{g}$ (green) and strong dissipation to $T_{2}=7T_{g}$ (red); the number of channel uses for the mixed-state reconstruction is not affected by the dissipation strength. The threshold below which SVD values of the output states are considered zero was taken to be $\epsilon=0.01$.
The advantage of the pure-state and mixed-state reconstruction methods compared to the Choi reconstruction results from exploiting the near-unitarity of the dynamical map. It is an advantage not only in terms of channel uses, as expressed in Eq. (28), but also in terms of ease of implementation, as our reconstruction methods do not require the preparation of entangled input states. They share this feature with low-rank process tomography methods [24] which suggests to compare to these protocols in terms of both implementation and scaling. A key difference of our approach is that no optimization, typically carried out via a semi-definite program [24], is required; rather Eqs. (4) and (5) provide explicit formulae for reconstructing the unitary directly from the output states.
In order to compare the scaling of the number $M$ of channel uses as a function of system size, we proceed in two steps. First, we determine $M$ for each method with Eq. (23) for different pure dephasing strengths. We only consider pure dephasing here as the worse case since amplitude damping may decrease the rank of output states $r_{\text{out}}$ and thus result in better scaling. Since $r_{\text{out}}$ is the determining factor for the minimum number of required channel uses $M$ in Eq. (23), we extract it numerically for noisy randomly chosen $N$-qubit operations. The resulting numbers of channel uses are shown in Fig. 7, confirming the scaling predictions of Eqs. (25) and (26): For dynamics close to unitarity, i.e. weak dissipation, the rank of the output states does not change substantially such that the number of channel uses of the pure state reconstruction is identical to the purely unitary case. Furthermore, Fig. 7 also demonstrates that the prediction of Eq. (28) holds: If the decoherence does not exceed a certain threshold, the pure state reconstruction requires less resources (channel uses) than the mixed state one.
In the second step, we can now compare these results to existing low-rank process tomography methods:
  1. 1.
    For a two-qubit process close to unitarity, the methods presented in Ref. [24] and Ref. [25] require between $M\sim 50$ and $M\sim 150$ channel uses, whereas Eq. (25) implies between $M=8^{2}=64$ (for $r\approx 1$) and $M=2^{N}8^{N}=256$ (for $r\approx\sqrt{d}$) channel uses for pure state reconstruction.
  2. 2.
    The reconstruction of a 3-qubit Toffoli gate under weak noise requires between $M\sim 320$ and $M\sim 1000$ channel uses depending on the strength of the noise [24]. In comparison, the scaling of our pure-state reconstruction of a three-qubit gate under weak noise requires $M=8^{3}=512$ channel uses.
These two examples show that our methods are competitive in terms of the required resources for small systems.
For larger systems, we consider the 7-qubit channel tomography of Ref. [26] reporting $M_{R}=10^{9}$ channel uses, compared to our value of $M=16^{7}$, and a unitary 10-qubit channel tomography [27] for which $M=10^{12}$ channel uses are used compared to $M=16^{10}$ channel uses with our method. We thus find that the protocols introduced here are competitive with state-of-the-art approaches for low-rank quantum processes.
@@ VI Conclusions
We have presented two methods to approximate the unitary part of a non-unitary evolution using a set of $d+1$ pure, resp. two mixed, input states, where $d$ is the dimension of Hilbert space. These states were designed to collect all information relevant to certify a quantum channel, i.e. to compare the realized quantum evolution to a desired unitary [5]. The present work has extended this approach to the reconstruction of an unknownunknown unitary, providing explicit formulae that require only knowledge of the output states. We have shown that reconstruction works well for no or weak decoherence. As expected, it struggles for strong decoherence where degeneracies in the image of the dynamical map hamper the meaningful identification of a specific unitary. We find that reconstruction using pure input states offers a good compromise between accuracy and resource cost. If accuracy is paramount, the Choi reconstruction provides the best results but requires a number of channel uses that quickly becomes unfeasible with increasing system size. When comparing to compressed sensing methods [21, 22, 4], in the regime of weak decoherence, the minimum number of channel uses is decreased from $M=\mathcal{O}(16^{N})$ to $M=\mathcal{O}(8^{N})$, where $N$ is the number of qubits, for the pure-state reconstruction. Moreover, comparing to compressed sensing methods exploiting the low rank of the quantum channel under weak noise [24, 25, 26, 27], we find the resource scaling of our protocols to be competitive for both small and large $N$. Our approach differs from these tomographic approaches which rely on a least-squares fit over sets of quantum channel [4]. Although this precludes to obtain analytical bounds using techniques from, e.g., compressed sensing, the lack of a numerical optimization removes a point of failure and reduces computational effort.
@@ Appendix B CNOT-like gates for multiple qubits
For $N=2$, we consider the CNOT gate where the first (second) qubit is the control (target) qubit. For multiple qubits, i.e. $N\in\{3,4,5\}$, we assume that the extra qubits (as compared to the $N=2$ case) are additional control qubits while the target qubit remains the second one. We also assume that only a single qubit, namely the target qubit dissipates, so that the absolute dissipation is the same for every system size. Figures 8 and 9 show the reconstruction error of the pure-state and mixed-state reconstruction methods respectively. They demonstrate that both methods can be used to reconstruct CNOT-like gates for multiple qubits, as indicated by the continuous reconstruction error graphs. However, in the left panel of Fig. 9 we notice that for the mixed-state reconstruction there are degeneracies in the image (see red graph), that do not appear in the pure-state reconstruction, cf. left panel of Fig. 8. When comparing the reconstruction errors of both methods, we again notice that the pure-state reconstruction performs better (compare position of the horizontal dotted blue line on the y-axis for the pure-state and mixed-state reconstruction).
Figure 8: Behavior of reconstruction error $1-F_{G}(U_{0},U_{RC})$ of the pure-state reconstruction method of CNOT-like gates for four designated $T_{1}$ times as a function of the number of qubits NN.
Figure 9: Behavior of reconstruction error $1-F_{G}(U_{0},U_{RC})$ of the mixed-state reconstruction method of CNOT-like gates for four designated $T_{1}$ times as a function of the number of qubits NN.